# FeDaL: Federated Dataset Learning for Time Series Foundation Models

### v1

@@ C.5 Federated Scaling Behaviors
While prior work explores scaling laws of centralized TSFMs in terms of model and data size, we instead investigate scaling behaviors in the federated setting, focusing on three key factors: (i) pretraining data size, (ii) number of clients, and (iii) client participation rate. Specifically, we vary: (i) Data size from $\{40\text{B},80\text{B},120\text{B},160\text{B},200\text{B},231\text{B}\}$ with a fixed client count of 174; (ii) Number of clients from $\{30,70,110,140,174\}$ using the full 231B dataset; (iii) Client participation rate from $\{10\%,30\%,50\%,70\%,100\%\}$ with the original pretraining setup. All experiments follow the training protocol in Section 4.3. Results are presented in Figure˜9. We observe that: (1) Increasing data size consistently improves downstream performance, even with a fixed client count; (2) More clients (with constant total data size) lead to better representations, suggesting improved learning of diverse, domain-specific patterns; (3) Higher client participation rates yield stronger results, likely due to more effective aggregation and reduced drift. These findings suggest that federated TSFM pretraining benefits from scaling in data and client dimensions, much like centralized pretraining. However, unlike the “scale-is-all” mindset focused on model size, our results highlight the importance of expanding dataset coverage and client participation as more efficient and federated-aligned strategies for improving generalization.

### v2

@@ C.7 Federated Scaling Behaviors
While prior work explores scaling laws of centralized TSFMs in terms of model and data size, we instead investigate scaling behaviors in the federated setting, focusing on three key factors: (i) pretraining data size, (ii) number of clients, and (iii) client participation rate. Specifically, we vary: (i) Data size from $\{40\text{B},80\text{B},120\text{B},160\text{B},200\text{B},231\text{B}\}$ with a fixed client count of 174; (ii) Number of clients from $\{30,70,110,140,174\}$ using the full 231B dataset; (iii) Client participation rate from $\{10\%,30\%,50\%,70\%,100\%\}$ with the original pretraining setup. All experiments follow the training protocol in Section 4.3. Results are presented in Figure 10. We observe that: (1) Increasing data size consistently improves downstream performance, even with a fixed client count; (2) More clients (with constant total data size) lead to better representations, suggesting improved learning of diverse, domain-specific patterns; (3) Higher client participation rates yield stronger results, likely due to more effective aggregation and reduced drift. These findings suggest that federated TSFM pretraining benefits from scaling in data and client dimensions, much like centralized pretraining. However, unlike the “scale-is-all” mindset focused on model size, our results highlight the importance of expanding dataset coverage and client participation as more efficient and federated-aligned strategies for improving generalization.
@@
C.8 Variance Estimates
All experiments (except the zero-shot forecasting setting) were conducted three times, and we report the averaged results. Variance estimates for representative configurations are now provided in Table 27-Table 30.
Table 27: Federated representation learning results with variance estimates. For USTD, only the H2 setting is reported as a representative case.
Method USTD CTSD
20% 35% 50% 75% 90% 20% 35% 50% 75% 90%
FedAvg 0.404 ($\pm$ 0.007) 0.489 ($\pm$ 0.011) 0.565 ($\pm$ 0.004) 0.602 ($\pm$ 0.007) 0.901 ($\pm$ 0.003) 0.350 ($\pm$ 0.002) 0.388 ($\pm$ 0.008) 0.405 ($\pm$ 0.011) 0.480 ($\pm$ 0.005) 0.652 ($\pm$ 0.005)
FedProx 0.401 ($\pm$ 0.004) 0.486 ($\pm$ 0.010) 0.553 ($\pm$ 0.013) 0.603 ($\pm$ 0.009) 0.889 ($\pm$ 0.011) 0.336 ($\pm$ 0.009) 0.390 ($\pm$ 0.004) 0.400 ($\pm$ 0.009) 0.454 ($\pm$ 0.019) 0.638 ($\pm$ 0.004)
FedPer 0.385 ($\pm$ 0.002) 0.578 ($\pm$ 0.008) 0.533 ($\pm$ 0.010) 0.580 ($\pm$ 0.007) 0.863 ($\pm$ 0.012) 0.330 ($\pm$ 0.001) 0.381 ($\pm$ 0.004) 0.395 ($\pm$ 0.004) 0.439 ($\pm$ 0.008) 0.606 ($\pm$ 0.015)
FedRep 0.413 ($\pm$ 0.007) 0.560 ($\pm$ 0.003) 0.537 ($\pm$ 0.005) 0.592 ($\pm$ 0.009) 0.860 ($\pm$ 0.007) 0.352 ($\pm$ 0.003) 0.373 ($\pm$ 0.003) 0.386 ($\pm$ 0.009) 0.440 ($\pm$ 0.010) 0.598 ($\pm$ 0.006)
FFTS 0.327 ($\pm$ 0.001) 0.410 ($\pm$ 0.010) 0.510 ($\pm$ 0.011) 0.584 ($\pm$ 0.007) 0.823 ($\pm$ 0.008) 0.300 ($\pm$ 0.002) 0.357 ($\pm$ 0.004) 0.386 ($\pm$ 0.002) 0.440 ($\pm$ 0.007) 0.598 ($\pm$ 0.005)
Stand-alone 0.380 ($\pm$ 0.002) 0.450 ($\pm$ 0.013) 0.521 ($\pm$ 0.001) 0.616 ($\pm$ 0.013) 0.870 ($\pm$ 0.006) 0.342 ($\pm$ 0.012) 0.381 ($\pm$ 0.004) 0.395 ($\pm$ 0.006) 0.470 ($\pm$ 0.014) 0.646 ($\pm$ 0.018)
FeDaL (Ours) 0.300 ($\pm$ 0.005) 0.422 ($\pm$ 0.003) 0.489 ($\pm$ 0.009) 0.549 ($\pm$ 0.005) 0.795 ($\pm$ 0.008) 0.310 ($\pm$ 0.007) 0.319 ($\pm$ 0.007) 0.343 ($\pm$ 0.002) 0.405 ($\pm$ 0.005) 0.560 ($\pm$ 0.011)
Table 28: Long-term forecasting and imputation Results with Variance Estimates. We use the forecasting horizon of 720 and masking ratio of 50% with MSE as the representative case.
Datasets FeDaL (Long-term Forecasting) FeDaL (Imputation)
ETTh1 0.457 ($\pm$ 0.003) 0.030 ($\pm$ 0.001)
ETTh2 0.401 ($\pm$ 0.004) 0.022 ($\pm$ 0.002)
ETTm1 0.464 ($\pm$ 0.004) 0.086 ($\pm$ 0.001)
ETTm2 0.397 ($\pm$ 0.002) 0.048 ($\pm$ 0.002)
Weather 0.359 ($\pm$ 0.005) 0.024 ($\pm$ 0.003)
Table 29: Short-term forecasting results with standard deviations (SMAPE Report).
Intervals Yearly Quarterly Monthly Others
FeDaL 13.102 ($\pm$ 0.11) 9.808 ($\pm$ 0.10) 12.124 ($\pm$ 0.19) 4.508 ($\pm$ 0.07)
Table 30: Anomaly detection results with standard deviations (F1-score Report).
Datasets SMD MSL SMAP SWaT PSM
FeDaL 88.46 ($\pm$ 0.26) 89.05 ($\pm$ 0.12) 71.70 ($\pm$ 0.09) 95.40 ($\pm$ 0.15) 98.88 ($\pm$ 0.21)
The Use of Large Language Models
In preparing this paper, we used a Large Language Model as a general-purpose writing assistant. Its role was limited to language refinement, such as improving grammar, sentence flow, and vocabulary usage to meet academic writing standards. All substantive aspects of the research (including conceptualization, experimental design, data analysis, and drawing conclusions) were carried out independently by the authors. The LLM was not employed for generating scientific content. We take full responsibility for the authenticity, originality, and integrity of this work.